OpenStudy (anonymous):

solve and express in a+bi form x^4 -7x^2-8

OpenStudy (anonymous):

\[(x^2+1)(x^2-8)=0\]\[x= +/- 2\sqrt{2} + 0i \]\[x= 0 +/- 1i\]

OpenStudy (anonymous):

-i

jhonyy9 (jhonyy9):

note x2=t and will get t2-7t-8=(t-8)(t+1) and from this result t_1=8 and t_2= -1 so than t_1=8 result x_1,_2= +/- 2sqrt2 +0i and than t_2= -1 --- x_3,_4= +/- sqrt(-1) = 0+/- i